# Steady state diagram of interacting fermionic atoms coupled to dissipative cavities

### v1

@@ VI Conclusions
To summarize, we have investigated the steady state phase diagrams of fermionic atoms in optical lattices coupled to an optical cavity at both low filling and half filling. We take the quantum fluctuations in the atoms-cavity coupling into account, going beyond the T=0T\!=\!0 mean-field method. By including the fluctuations within the many-body adiabatic elimination method we arrive at mixed steady states described by an effective temperature. The finite value of the effective temperature shifts the transition between the normal and self-organized phases with respect to the ground-state (T=0T\!=\!0) approximation. We substantiate our results around the phase transition with approximate analytical expressions in various limits. Deep in the self-organized phase, a saturation in the cavity field occupation is observed and by employing an analytical approximation we show that the photon number becomes independent of the atoms-cavity coupling strength.
Our approach can capture the dynamical energy exchange between the atoms and the cavity field mediated by the fluctuations. Thus, we identified cavity-cooling processes as well as many-body-cooling processes and associated them with resonances between the photon energy and gap in the effective atomic model. In the vicinity of these resonances, the energy transfer between excited eigenstates of the atomic system to the photonic mode is facilitated. At low filling, as exemplarily shown with results at quarter filling, we find the occurrence of a fluctuation-induced bistability region Tolle et al. (2025a), which does not exist in the T=0T\!=\!0 MF solution. The emergence of the bistability relies on the many-body cooling processes and can only be capture due to the self-consistent determination of the effective temperature present in the many-body adiabatic elimination approach.

### v2

@@ VI Conclusions
To summarize, we have investigated the steady state phase diagrams of fermionic atoms in optical lattices coupled to an optical cavity at both low filling and half filling. We take the quantum fluctuations in the atoms-cavity coupling into account, going beyond the T=0T\!=\!0 mean-field method. By including the fluctuations within the many-body adiabatic elimination method we obtain mixed steady states described by an effective temperature. We investigate the self-organization transition, where the finite value of the effective temperature shifts the transition between the normal and self-organized phases with respect to the ground-state (T=0T\!=\!0) approximation. We analyze both the low filling scenario, in which the transition is continuous, and the commensurate half-filled case where a bistability occurs at the phase transition. We show that in the limit of strong dissipation the quantum fluctuations can reduce the width of the bistability arising at half-filling with respect to the predictions of the T=0T\!=\!0 mean-field method. We substantiate our results around the phase transition with approximate analytical expressions in various limits. Deep in the self-organized phase, a saturation in the cavity field occupation is observed and by employing an analytical approximation we show that the photon number becomes independent of the atoms-cavity coupling strength.
Our approach can capture the dynamical energy exchange between the atoms and the cavity field mediated by the fluctuations. Thus, we identified cavity-cooling processes as well as many-body cooling processes and associated them with resonances between the photon energy and gap in the effective atomic model. In the vicinity of these resonances, the energy transfer between excited eigenstates of the atomic system to the photonic mode is facilitated. At low filling, as exemplarily shown with results at quarter filling, we find the occurrence of a fluctuation-induced bistability region Tolle et al. (2025a), which does not exist in the T=0T\!=\!0 MF solution. The emergence of the bistability relies on the many-body cooling processes and can only be capture due to the self-consistent determination of the effective temperature present in the many-body adiabatic elimination approach.